OpenStudy (anonymous):

How far does a boat travel in 5 hours at 32 miles per hour?

OpenStudy (anonymous):

5*32= 160 miles.....

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

distance=speed*time distance=5*32 distance=160 miles